find the locus of a point whose distance from the point 1, 2 is equal to its distance from the axis of y

Grade:11

1 Answers

Arun
6 years ago
Dear kamal
Let the point (x, y)
Now
Its distance from y -axis = x
Hence
\large \sqrt(x-1)² + (y-2)² = x
On squaring both the sides
x² +1 -2x + y² + 4 -4y = x²
y² -4y +4 = 2x-1
(y-2)² = 4*(1/2) (x -1/2)
 
It is a parabola
 
Regards
Arun (askIITians forum expert)